OpenStudy (anonymous):

Use the substitution method to solve the system of equations. x + 2y = 15 2x – 3y = 2

Parth (parthkohli):

\[\begin{align}x &=&15 - 2y \\ 2(15 - 2y) - 3y & = & 2\end{align}\]

OpenStudy (anonymous):

how did you come up with that??

OpenStudy (anonymous):

x+2y=15 or, x=15-2y...............(i) Now, 2x-3y=2 or,2(15-2y)=2 (using the value of x from equation (i) ) or,30-4y=2 or,-4y=2-30 or,-4y=-28 or, y=-28/-4 So, y=3 then put this value y=3 in equation (i). x=15-2(3) or, x=15-6 So, x=9 hence, the required values are x=9 and y=3